OpenStudy (mlgpro):

Two trees stand side-by-side. The shorter tree is four feet tall. Its shadow is 9 feet long. The taller tree is eighteen feet tall. How long is its shadow? 4.5 feet 8 feet 23 feet 40.5 feet

OpenStudy (mlgpro):

@JuniorEinstein1 this has 2 answers

OpenStudy (anonymous):

hese will form two similar triangles so ratios of 2 heights = ratio of the 2 shadows 4 / 18 = 9 / x solve for x

OpenStudy (anonymous):

cross multiply: 4x = 9*18 4x = 162 can you continue?

OpenStudy (mlgpro):

the 2 formulas are x+4=9 x=5 x*4=9 x=2.25

OpenStudy (anonymous):

i just told what u need to do

OpenStudy (mlgpro):

i didnt see sorry

OpenStudy (anonymous):

im not gunna give you the answer becasue CoC wont allow me

OpenStudy (mlgpro):

i do it

OpenStudy (mayankdevnani):

 tan theta is constant. so use tan theta \[\huge \bf \tan \theta=\frac{4}{9}=\frac{18}{x}\] so solve for x